A firm is must choose to buy the GSU-3300 or the UGA-3000. Both machines make the firm’s production process more efficient which in turn increases incremental cash flows. The GSU-3300 produces incremental cash flows of $24,455.00 per year for 8 years and costs $98,558.00. The UGA-3000 produces incremental cash flows of $27,924.00 per year for 9 years and cost $126,561.00. The firm’s WACC is 7.03%. What is the equivalent annual annuity of the UGA-3000?
Answer #1

The first step is to calculate the present value of cash flows of UGA-3000:

PV of cash flows = Incremental cash flows * (1 - (1+R)^(-n)) / R

where R = 7.03% and N = 9

= 27924 * (1 - (1+7.03%)^-9) / 0.0703

= 181699.39

NPV = PV of cash flows - Initial investment = 181699.39 - 126561

= 55138.39

Equivalent annual annuity = NPV * R / (1 - (1+R)^(-n))

= 55138.39 * 7.03% / (1 - (1+7.03%)^(-9))

= 8473.80